OpenStudy (anonymous):

ab, for a =60 and b=6 5/18

OpenStudy (anonymous):

turn the mixed number 6 and 5/18 into a single fraction: 18*6 + 5 = 113 so 113/18...

OpenStudy (anonymous):

then 60 is the same as 60/1, so : 60/1 * 113/18 cancel first, then multiply across the numerator and denominator

OpenStudy (anonymous):

6 5/18 = 113/18 now: a*b=60*113/18=20*113/6=10*113/3=1130/3